# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Abstract
Models based on recursive adaptive partitioning such as decision trees and their ensembles are popular for high-dimensional regression as they can potentially avoid the curse of dimensionality. Because empirical risk minimization (ERM) is computationally infeasible, these models are typically trained using greedy algorithms. Although effective in many cases, these algorithms have been empirically observed to get stuck at local optima. We explore this phenomenon in the context of learning sparse regression functions over $d$ binary features, showing that when the true regression function $f^{*}$ does not satisfy Abbe et al. (2022)’s Merged Staircase Property (MSP), greedy training requires $\exp(\Omega(d))$ to achieve low estimation error. Conversely, when $f^{*}$ does satisfy MSP, greedy training can attain small estimation error with only $O(\log d)$ samples. This dichotomy mirrors that of two-layer neural networks trained with stochastic gradient descent (SGD) in the mean-field regime, thereby establishing a head-to-head comparison between SGD-trained neural networks and greedy recursive partitioning estimators. Furthermore, ERM-trained recursive partitioning estimators achieve low estimation error with $O(\log d)$ samples irrespective of whether $f^{*}$ satisfies MSP, thereby demonstrating a statistical-computational trade-off for greedy training. Our proofs are based on a novel interpretation of greedy recursive partitioning using stochastic process theory and a coupling technique that may be of independent interest.
@@ 1 Introduction
Here, the covariate space $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$, $\mathbf{X}_{1},\mathbf{X}_{2},\ldots\mathbf{X}_{n}$ are drawn i.i.d. from a distribution $\nu$ on $\mathcal{X}$, $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are noise variables drawn i.i.d. from a zero-mean noise distribution on $\mathbb{R}$, with $\varepsilon_{i}\perp\!\!\!\perp\mathbf{X}_{i}\eqqcolon(X_{i1},X_{i2},\ldots,X_{id})$, and $f^{*}\colon\mathcal{X}\to\mathbb{R}$ is the conditional expectation function of $Y$ given $\mathbf{X}=\mathbf{x}$. The observed data is denoted as $\mathcal{D}_{n}\coloneqq\left\{(\mathbf{X}_{i},Y_{i})\colon i=1,\ldots,n\right\}$. We will compare and contrast various tree-based estimators for $f^{*}$ , evaluating the accuracy of an estimate $\hat{f}(-;\mathcal{D}_{n})$ using the $L^{2}$ risk (or estimation error)
@@
and the accuracy of an estimator $\hat{f}(-;-)$ using the expected $L^{2}$ risk, $\mathbb{E}_{\mathcal{D}_{n}}\left\{R\left(\hat{f}(-;\mathcal{D}_{n}),f^{*}\right)\right\}$. Note that throughout this paper, we will use the convention that random variables are denoted in upper case while any fixed value they may attain is denoted in lower case. Furthermore, vectors will be denoted using boldface, whereas scalars will be denoted using regular font. We denote $[d]\coloneqq\left\{1,2,\ldots,d\right\}$. To supplement standard Big-$O$ notation, we use $\tilde{O}$ to suppress poly-logarithmic factors.
1.1 Consistency of ERM trees, Greedy Trees, and Non-Adaptive Trees
Machine learning models are often fit by solving empirical risk minimization (ERM). Solutions to ERM over appropriate classes of decision tree functions can be shown to be consistent, but unfortunately are computationally infeasible to obtain. For instance, ERM for classification trees is known to be NP-hard in the worst case (Hyafil and Rivest, 1976).
As a result, practitioners make use of greedy algorithms such as CART  (Breiman et al., 1984), ID3 (Quinlan, 1986), or C4.5 (Quinlan, 1993). Such algorithms grow a tree in a top-down manner, with the split for a given node chosen by optimizing a local node-specific objective. Once made, a split is never reversed. [^1] Decision trees are sometimes pruned as a one-off post-processing procedure after they are fully grown. These heuristics seem to work well in practice—both RFs and gradient boosting are built out of CART trees—but are notoriously difficult to analyze. Indeed, early efforts to prove consistency of greedy regression and classification trees under standard smoothness conditions on $f^{*}$ had to further assume that the bias of the tree converged to zero (Breiman et al., 1984; Chaudhuri et al., 1995). More recent works were able to replace this assumption with more reasonable alternatives on $f^{*}$ such as additivity  (Scornet et al., 2015; Klusowski, 2020; Klusowski and Tian, 2024) or a submodularity condition (Syrgkanis and Zampetakis, 2020). Chi et al. (2022) further introduced and derived consistency under the sufficient impurity decrease (SID) condition, which we will define more formally later in this paper. Mazumder and Wang (2024) later improved upon the consistency rates of Chi et al. (2022) and showed that the SID condition is satisfied for larger classes of models.
In another line of work, researchers studied and were able to prove consistency for decision trees making splits that are non-adaptive (i.e., independent of the response). These algorithms include trees that make completely random splits (Geurts et al., 2006; Biau et al., 2008; Biau, 2012; Wager and Athey, 2018; Klusowski, 2021), Mondrian trees which make splits according to a Bayesian modeling of the covariate distribution (Roy and Teh, 2008; Lakshminarayanan et al., 2014; Mourtada et al., 2017; Cattaneo et al., 2023) and $k$-Potential Nearest Neighbors (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet, 2016; Biau and Scornet, 2016; Shi et al., 2024). Despite possessing theoretical guarantees requiring fewer assumptions, such algorithms exhibit performance gaps compared to greedy trees, in particular, in terms of adaptivity to heterogeneous truth.
@@ 1.2 High-Dimensional Consistency
From the discussion thus far, we gather that considering nonparametric consistency alone does not explain why greedy trees have been so successful relative to non-adaptive trees, let alone other machine learning algorithms such as $k$-nearest neighbors or support vector machines. To explain this, we compare the performance of these algorithms in a high-dimensional setting ($n,d\to\infty$) with sparsity constraints on $f^{*}$, i.e., if $f^{*}$ only depends on a constant number of $s$ covariates. Informally, we say that an estimator avoids the curse of dimensionality, or that it is high-dimensional consistent, if its expected $L^{2}$ risk converges to zero under a scaling regime satisfying $\log n=o(d)$.
We first argue that non-adaptive trees are not high-dimensional consistent [^2] See Appendix A for a formal definition of non-adaptive tree models and a rigorous proof of this statement.: Assuming relatively balanced splits, the average depth of a tree is $O(\log n)$. Hence, $\log n=o(d)$ implies that not every covariate can receive a split, or in other words, that the tree needs to be able to selectively split on relevant covariates, which contradicts the definition of non-adaptivity. In contrast, several works have showed that CART and RFs are high-dimensional consistent given various condition on $f^{*}$ (e.g., Syrgkanis and Zampetakis (2020) for binary covariates and assuming submodularity, Klusowski and Tian (2024) for arbitrary covariates and assuming additivity, Chi et al. (2022) for arbitrary covariates and assuming SID). This helps to explain the successful application of RFs to high-dimensional problems such as genomics in which $n\ll d$.
@@
We do this in the context of binary covariates under the uniform distribution, which although unlikely to describe any real dataset, provides a useful setting where both results and calculations can be stated cleanly and precisely, thereby illustrating the essence of the strengths and limitations of using a greedy splitting rule. We will later also provide extensions of these results to more general settings.
@@
for some nonnegative integer $r$, subsets $S_{1},S_{2},\ldots,S_{r}$ of $\left\{1,2,\ldots,d\right\}$ and nonzero coefficients $\alpha_{S_{j}},j=1,2,\ldots,d$. We say that $f^{*}$ satisfies the merged-staircase property (MSP) (Abbe et al., 2022) if the subsets $S_{1},S_{2},\ldots,S_{r}$ can be reordered such that for any $i\in\{1,2,\ldots,r\}$, we have
$\left|S_{i}\backslash\cup_{j=1}^{i-1}S_{j}\right|\leq 1.$ (5)
Note that any monomial of degree $k>1$ does not satisfy the MSP. In particular, the XOR function does not satisfy the MSP. On the other hand, a positive example is provided by the function
@@
Here, genericity of $\left\{\alpha_{S_{i}}\right\}_{i=1}^{r}$ is satisfied with probability 1 whenever they are drawn from a distribution with a density. Up to this asterisk, we see that high-dimensional consistency for CART can be completely characterized by MSP, a combinatorial condition on Fourier coefficients. In addition, we see in fact that CART behaves similarly to ERM trees when MSP holds, and similarly to non-adaptive trees when it does not, thereby also establishing performance gaps: In the former case, CART sharply improves upon non-adaptive trees, and in the latter case, it performs much worse compared to ERM. Finally, we remark that our lower bounds when $f^{*}$ does not satisfy MSP also hold more broadly for RFs, as well as for trees and ensembles grown with other greedy recursive partitioning strategies. [^5] It is possible to extend the upper bound to various forms of RFs, but we leave this to future work.
@@ 1.4 Marginal Signal Bottleneck
On the other hand, it is easy to show that MSP together with genericity implies the SID condition. In the setting of uniform binary features, this condition asserts a lower bound for what we call the marginal signal,
@@ 1.7 Comparisons with Neural Networks Trained by SGD
In computational learning theory, the problem of “learning” (i.e., achieving small estimation error for) sparse Boolean monomials, also called parities, from i.i.d. noiseless observations (i.e., $\varepsilon_{i}=0$ in (1)) is known to be statistically easy but computationally hard (Barak et al., 2022), and hence has served as a useful benchmark for learning algorithms and computational frameworks. Recently, this has become a benchmark for studying neural networks (NNs) trained using SGD, which combined with our results, allows us to perform, to our knowledge, the first theoretical head-to-head comparison between this class of algorithms and greedy regression trees and ensembles.
@@
Of further interest is Abbe et al. (2023)’s conjecture characterizing the sample complexity required to learn any given Boolean function in terms of its “leap complexity” (see Appendix J). If this conjecture is true, NNs trained with SGD outperform greedy regression trees and ensembles on all non-MSP  functions but perform more poorly in comparison on MSP functions. We discuss these comparisons more elaborately in Appendix J.
@@ 2.1 Cells, Partitions, and Splits
A cell $\mathcal{C}$ is a subcube of {±$\{\pm 1\}^{d}$ and is defined as
@@
for a subset of covariate indices $J(\mathcal{C})\subset[d]$ and a sequence of signs $z_{j}\in\left\{\pm 1\right\}$, $j\in J(\mathcal{C})$. Given a choice of covariate index $k\notin J(\mathcal{C})$ and sign $z\in\left\{\pm 1\right\}$, we define the split operator as producing the output $\mathcal{T}_{k,z}(\mathcal{C})\coloneqq\left\{\mathbf{x}\in\mathcal{C}\colon x_{k}=z\right\}$. $\mathcal{T}_{k,1}(\mathcal{C})$ and $\mathcal{T}_{k,-1}(\mathcal{C})$ are called the right and left child of cell $\mathcal{C}$. A partition $\mathcal{P}$ is a collection of disjoint cells whose union comprises {±$\{\pm 1\}^{d}$. The size of $\mathcal{P}$ is the number of cells it contains and is denoted $|\mathcal{P}|$. We say that another partition $\mathcal{P}^{\prime}$ is a refinement of $\mathcal{P}$, denoted $\mathcal{P}\subset\mathcal{P}^{\prime}$, if for any $\mathcal{C}^{\prime}\in\mathcal{P}^{\prime}$, there exists $\mathcal{C}\in\mathcal{P}$ such that $\mathcal{C}^{\prime}\subset\mathcal{C}$. We say that $\mathcal{P}^{\prime}$ is a simple refinement of $\mathcal{P}$, denoted $\mathcal{P}\subset_{s}\mathcal{P}^{\prime}$ if $\mathcal{P}\backslash\mathcal{P}^{\prime}=\{\mathcal{C}\}$, $\mathcal{P}^{\prime}\backslash\mathcal{P}=\{\mathcal{C}_{L}^{\prime},\mathcal{C}_{R}^{\prime}\}$ and $\mathcal{C}=\mathcal{C}_{L}^{\prime}\cup\mathcal{C}_{R}^{\prime}$. In other words, $\mathcal{P}^{\prime}$ is obtained from $\mathcal{P}$ by applying the split operator to $\mathcal{C}$. We use $I(\mathcal{C})$ to denote the indices of observations whose covariate vectors lie in $\mathcal{C}$, i.e., $I(\mathcal{C})\coloneqq\left\{i\in[n]\colon\mathbf{x}_{i}\in\mathcal{C}\right\}$, and let $N(\mathcal{C})\coloneqq\left|I(\mathcal{C})\right|$. Furthermore, we denote the mean response in $\mathcal{C}$ as $\bar{Y}_{\mathcal{C}}\coloneqq\frac{1}{N(\mathcal{C})}\sum_{i\in I(\mathcal{C})}Y_{i}$.
@@ 2.2 Regression Tree Models
Note that every piecewise constant function $g$ on {±$\{\pm 1\}^{d}$ can be parameterized using a partition $\mathcal{P}$ with a fixed ordering of its constituent cells $\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{|\mathcal{P}|}$, and a set of labels $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\ldots,\mu_{|\mathcal{P}|})$ representing the values taken by $g$ on each cell. In other words, we have
@@
A binary, axis-aligned regression tree model is a special type of piecewise constant function, one whose partition $\mathcal{P}$ arises from recursive applications of split operators $\mathcal{T}_{\cdot,\cdot}$ to {±$\{\pm 1\}^{d}$ and the resulting cells. More precisely, there exists a sequence of simple refinements, {{±$\{\{\pm 1\}^{d}\}=\mathcal{P}_{1}\subset_{s}\mathcal{P}_{1}\subset_{s}\cdots\subset_{s}\mathcal{P}_{|\mathcal{P}|}=\mathcal{P}$. This recursive structure can be represented as a labeled binary tree $\mathfrak{T}$ in which each node corresponds to a cell obtained in the recursion and parent-child relationships correspond to applications of $\mathcal{T}_{\cdot,\cdot}$. The root of $\mathfrak{T}$ is the entire covariate space {±$\{\pm 1\}^{d}$. As is common in the literature, we will abuse terminology and sometimes use the term “node” to refer to its corresponding cell. Each internal node is labeled with the covariate used to split it, called its split covariate and denoted as $k(\mathcal{C})$. The non-internal nodes $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{|\mathcal{P}|}$ are called leaves, have empty labels, and comprise the final partition $\mathcal{P}=\mathcal{P}(\mathfrak{T})$. For convenience, we also slightly abuse notation, writing $\left|\mathfrak{T}\right|=\left|\mathcal{P}(\mathfrak{T})\right|$ and $g(-;\mathfrak{T},\boldsymbol{\mu})=g(-;\mathcal{P}(\mathfrak{T}),\boldsymbol{\mu})$. We call $\mathfrak{T}$ the tree structure of the model and $\boldsymbol{\mu}$ its leaf labels. The depth of $\mathfrak{T}$ is the maximum length of a root-to-leaf path, that is, a sequence of nodes $\mathcal{C}_{0}\supset\mathcal{C}_{1}\supset\cdots\supset\mathcal{C}_{l}$, where $\mathcal{C}_{j}$ is a parent of $\mathcal{C}_{j+1}$ in $\mathfrak{T}$ for $j=0,1,\ldots,{l-1}$, $\mathcal{C}_{0}$ is the root and $\mathcal{C}_{l}$ is a leaf node.

### v3

@@ Abstract
Models based on recursive adaptive partitioning such as decision trees and their ensembles are popular for high-dimensional regression as they can potentially avoid the curse of dimensionality. Because empirical risk minimization (ERM) is computationally infeasible, these models are typically trained using greedy algorithms. Although effective in many cases, these algorithms have been empirically observed to get stuck at local optima. We explore this phenomenon in the context of learning sparse regression functions over $d$ binary features, showing that when the true regression function $f^{*}$ does not satisfy Abbe et al. (2022)’s Merged Staircase Property (MSP), a form of heredity restriction similar to that used in classical ANOVA modeling, greedy training requires $\exp(\Omega(d))$ samples to achieve low estimation error. Conversely, when $f^{*}$ does satisfy MSP, greedy training can attain small estimation error with only $O(\log d)$ samples. This dichotomy mirrors that of two-layer neural networks trained with stochastic gradient descent (SGD) in the mean-field regime, thereby establishing a head-to-head comparison between SGD-trained neural networks and greedy recursive partitioning estimators. Furthermore, ERM-trained recursive partitioning estimators achieve low estimation error with $O(\log d)$ samples irrespective of whether $f^{*}$ satisfies MSP, thereby demonstrating a statistical-computational trade-off for greedy training. Our proofs are based on a novel interpretation of greedy recursive partitioning using stochastic process theory and a coupling technique that may be of independent interest.
@@ 1 Introduction
Here, the covariate space $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$, $\mathbf{X}_{1},\mathbf{X}_{2},\ldots\mathbf{X}_{n}$ are drawn i.i.d. from a distribution $\nu$ on $\mathcal{X}$, $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are noise variables drawn i.i.d. from a zero-mean noise distribution on $\mathbb{R}$, with $\varepsilon_{i}\perp\!\!\!\perp\mathbf{X}_{i}\eqqcolon(X_{i1},X_{i2},\ldots,X_{id})$, and $f^{*}\colon\mathcal{X}\to\mathbb{R}$ is the conditional expectation function of the response $Y$ given $\mathbf{X}=\mathbf{x}$. The observed data is denoted as $\mathcal{D}_{n}\coloneqq\left\{(\mathbf{X}_{i},Y_{i})\colon i=1,\ldots,n\right\}$. We will compare and contrast various tree-based estimators for $f^{*}$ , evaluating the accuracy of an estimate $\hat{f}(-;\mathcal{D}_{n})$ using the $L^{2}$ risk (or estimation error)
@@
[…]
@@ 1.1 High-Dimensional Consistency
To establish statistical-computational trade-offs for regression trees, we analyze their performance in a high-dimensional setting ($n,d\to\infty$) with sparsity constraints on $f^{*}$, i.e., if $f^{*}$ only depends on a constant number of $s$ covariates. Informally, we say that an estimator avoids the curse of dimensionality, or that it is high-dimensional consistent, if its expected $L^{2}$ risk converges to zero under a scaling regime satisfying $\log n=o(d)$. It is easy to see that high-dimensional consistency requires feature selection: Assuming relatively balanced splits, the average depth of a tree is $O(\log n)$. Hence, $\log n=o(d)$ implies that not every covariate can receive a split. As such, totally randomized trees, as well as other such algorithms that grow decision trees independently of the observed responses, are not high-dimensional consistent.
Several works have showed that CART and RFs are high-dimensional consistent given various condition on $f^{*}$ (e.g., Syrgkanis and Zampetakis (2020) for binary covariates and assuming submodularity, Klusowski and Tian (2024) for arbitrary covariates and assuming additivity, Chi et al. (2022) for arbitrary covariates and assuming a condition called sufficient impurity decrease, which we will define more formally later in this paper). This helps to explain the successful application of RFs to high-dimensional problems such as genomics in which $n\ll d$.
@@
We do this in the context of binary covariates under the uniform distribution, which although unlikely to describe any real dataset, provides a useful setting where both results and calculations can be stated cleanly and precisely, thereby illustrating the essence of the strengths and limitations of using a greedy splitting rule.
@@
for some nonnegative integer $r$, subsets $S_{1},S_{2},\ldots,S_{r}$ of $\left\{1,2,\ldots,d\right\}$ and nonzero coefficients $\alpha_{S_{j}},j=1,2,\ldots,r$. We say that $f^{*}$ satisfies the Merged-Staircase Property (MSP) (Abbe et al., 2022) if the subsets $S_{1},S_{2},\ldots,S_{r}$ can be reordered such that for any $i\in\{1,2,\ldots,r\}$, we have
[…]
Note that any monomial of degree $k>1$ does not satisfy the MSP. In particular, the XOR function does not satisfy the MSP. On the other hand, a positive example satisfying MSP is provided by the function
@@
For further intuition, note that (4) can be interpreted as an ANOVA decomposition for $f^{*}(\mathbf{X})$ (Wahba, 1990). Each term $\chi_{S_{i}}$ is a contrast that captures either a main effect (if $S_{i}$ is a singleton) or a $k$-way interaction (if $S_{i}$ has cardinality $k>1$). Meanwhile, the coefficient $\alpha_{S_{i}}$ represent the effect size. The MSP imposes a type of heredity [^3] The MSP is related but not equivalent to the conditions of weak and strong heredity often enforced by ANOVA modeling in practice (Wu and Hamada, 2011). Weak heredity, which asserts that an interaction is present only if at least one of its lower-order parents is present, is equivalent to the Staircase Property formulated by Abbe et al. (2021a) to analyze neural networks, and is therefore stronger than the MSP. Abbe et al. (2021a) later generalized the definition to the MSP, which turns out to offer a better characterization of learnability. constraint on the pattern of interactions present in the ANOVA model: We can order the nonzero effects so that each newly-introduced effect brings in at most one new variable that has not appeared in any previously included effect. In other words, interactions cannot appear “out of nowhere” combining many fresh variables at once.
@@
The formal statements of these results are provided in the forthcoming Theorems 5.1, 7.5, and 8.1, together with Propositions 6.1 and 6.2. Our upper bound is attained when using a minimum impurity decrease stopping condition (see (14) for details) with a fixed threshold $\gamma$ that is chosen above the noise level but below the signal strength. The lower bound does not depend on $\gamma$ or on the particular stopping/pruning choices.
Here, genericity of $\left\{\alpha_{S_{i}}\right\}_{i=1}^{r}$ holds with probability one whenever they are drawn from a distribution with a density (see the discussion in Section 7.1). Up to this asterisk, we see that high-dimensional consistency for CART can be completely characterized by MSP, a combinatorial condition on Fourier coefficients. In addition, we see in fact that CART behaves similarly to ERM trees when MSP holds, and similarly to non-adaptive trees when it does not, thereby also establishing performance gaps: In the former case, CART sharply improves upon non-adaptive trees, and in the latter case, it performs much worse compared to ERM. Finally, we remark that our lower bounds when $f^{*}$ does not satisfy MSP also hold more broadly for RFs, as well as for trees and ensembles grown with other greedy recursive partitioning strategies. [^4] It is possible to extend the upper bound to various forms of RFs, but we leave this to future work.
@@ 1.3 Marginal Signal Bottleneck
On the other hand, it is easy to show that MSP together with genericity implies Chi et al. (2022)’s sufficient impurity decrease (SID) condition. In the setting of uniform binary features, the SID condition asserts a lower bound for what we call the marginal signal,
@@ 1.6 Comparisons with Neural Networks Trained by SGD
We now compare and connect our results with the literature on training neural networks. Readers who are only interested in decision trees and ensembles can safely skip this sub-section.
In computational learning theory, the problem of “learning” (i.e., achieving small estimation error for) sparse Boolean monomials, also called parities, from i.i.d. noiseless observations (i.e., $\varepsilon_{i}=0$ in (1)) is known to be statistically easy but computationally hard (Barak et al., 2022), and hence has served as a useful benchmark for learning algorithms and computational frameworks. Recently, this has become a benchmark for studying neural networks (NNs) trained using SGD, which, when combined with our results, allows us to perform, to our knowledge, the first theoretical head-to-head comparison between this class of algorithms and greedy regression trees and ensembles.
@@
Of further interest is Abbe et al. (2023)’s conjecture characterizing the sample complexity required to learn any given Boolean function in terms of its “leap complexity”. If this conjecture is true, NNs trained with SGD outperform greedy regression trees and ensembles on all non-MSP  functions but perform more poorly in comparison on MSP functions. We discuss these comparisons more elaborately in Appendix K.
@@ 2.1 Cells, Partitions, and Splits
A cell $\mathcal{C}$ is a subcube of $\{\pm 1\}^{d}$ and is defined as
@@
for a subset of covariate indices $J(\mathcal{C})\subset[d]$ and a sequence of signs $z_{j}\in\left\{\pm 1\right\}$, $j\in J(\mathcal{C})$. Given a choice of covariate index $k\notin J(\mathcal{C})$ and sign $z\in\left\{\pm 1\right\}$, we define the split operator as producing the output $\mathcal{T}_{k,z}(\mathcal{C})\coloneqq\left\{\mathbf{x}\in\mathcal{C}\colon x_{k}=z\right\}$. $\mathcal{T}_{k,1}(\mathcal{C})$ and $\mathcal{T}_{k,-1}(\mathcal{C})$ are called the right and left child of cell $\mathcal{C}$. A partition $\mathcal{P}$ is a collection of disjoint cells whose union comprises $\{\pm 1\}^{d}$. The size of $\mathcal{P}$ is the number of cells it contains and is denoted $|\mathcal{P}|$. We say that another partition $\mathcal{P}^{\prime}$ is a refinement of $\mathcal{P}$, denoted $\mathcal{P}\subset\mathcal{P}^{\prime}$, if for any $\mathcal{C}^{\prime}\in\mathcal{P}^{\prime}$, there exists $\mathcal{C}\in\mathcal{P}$ such that $\mathcal{C}^{\prime}\subset\mathcal{C}$. We say that $\mathcal{P}^{\prime}$ is a simple refinement of $\mathcal{P}$, denoted $\mathcal{P}\subset_{s}\mathcal{P}^{\prime}$ if $\mathcal{P}\backslash\mathcal{P}^{\prime}=\{\mathcal{C}\}$, $\mathcal{P}^{\prime}\backslash\mathcal{P}=\{\mathcal{C}_{L}^{\prime},\mathcal{C}_{R}^{\prime}\}$ and $\mathcal{C}=\mathcal{C}_{L}^{\prime}\cup\mathcal{C}_{R}^{\prime}$. In other words, $\mathcal{P}^{\prime}$ is obtained from $\mathcal{P}$ by applying the split operator to $\mathcal{C}$. We use $I(\mathcal{C})$ to denote the indices of observations whose covariate vectors lie in $\mathcal{C}$, i.e., $I(\mathcal{C})\coloneqq\left\{i\in[n]\colon\mathbf{x}_{i}\in\mathcal{C}\right\}$, and let $N(\mathcal{C})\coloneqq\left|I(\mathcal{C})\right|$. Furthermore, we denote the mean response in $\mathcal{C}$ as $\bar{Y}_{\mathcal{C}}\coloneqq\frac{1}{N(\mathcal{C})}\sum_{i\in I(\mathcal{C})}Y_{i}$.
@@ 2.2 Regression Tree Models
Note that every piecewise constant function $g$ on $\{\pm 1\}^{d}$ can be parameterized using a partition $\mathcal{P}$ with a fixed ordering of its constituent cells $\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{|\mathcal{P}|}$, and a set of labels $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\ldots,\mu_{|\mathcal{P}|})$ representing the values taken by $g$ on each cell. In other words, we have
@@
A binary, axis-aligned regression tree model is a special type of piecewise constant function, one whose partition $\mathcal{P}$ arises from recursive applications of split operators $\mathcal{T}_{\cdot,\cdot}$ to $\{\pm 1\}^{d}$ and the resulting cells. More precisely, there exists a sequence of simple refinements, $\{\{\pm 1\}^{d}\}=\mathcal{P}_{1}\subset_{s}\mathcal{P}_{2}\subset_{s}\cdots\subset_{s}\mathcal{P}_{|\mathcal{P}|}=\mathcal{P}$. This recursive structure can be represented as a labeled binary tree $\mathfrak{T}$ in which each node corresponds to a cell obtained in the recursion and parent-child relationships correspond to applications of $\mathcal{T}_{\cdot,\cdot}$. The root of $\mathfrak{T}$ is the entire covariate space $\{\pm 1\}^{d}$. As is common in the literature, we will abuse terminology and sometimes use the term “node” to refer to its corresponding cell. Each internal node is labeled with the covariate used to split it, called its split covariate and denoted as $k(\mathcal{C})$. The non-internal nodes $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{|\mathcal{P}|}$ are called leaves, have empty labels, and comprise the final partition $\mathcal{P}=\mathcal{P}(\mathfrak{T})$. For convenience, we also slightly abuse notation, writing $\left|\mathfrak{T}\right|=\left|\mathcal{P}(\mathfrak{T})\right|$ and $g(-;\mathfrak{T},\boldsymbol{\mu})=g(-;\mathcal{P}(\mathfrak{T}),\boldsymbol{\mu})$. We call $\mathfrak{T}$ the tree structure of the model and $\boldsymbol{\mu}$ its leaf labels. The depth of $\mathfrak{T}$ is the maximum length of a root-to-leaf path, that is, a sequence of nodes $\mathcal{C}_{0}\supset\mathcal{C}_{1}\supset\cdots\supset\mathcal{C}_{l}$, where $\mathcal{C}_{j}$ is a parent of $\mathcal{C}_{j+1}$ in $\mathfrak{T}$ for $j=0,1,\ldots,{l-1}$, $\mathcal{C}_{0}$ is the root and $\mathcal{C}_{l}$ is a leaf node.